When 250 kg of copper (I) oxide is heated with 129 kg of copper (I) sulfide, 285 kg of copper is recovered. Calculate the theoretical yield of copper.
1
Expert's answer
2019-03-18T02:20:58-0400

2Cu2O + Cu2S -> 6Cu + SO2

n= m/M

n(Cu2O)= 250000g/143.09g/mol= 1747 mol

n(Cu2S)= 129000g/ 159.16 g/mol= 810.5 mol

Find limiting reactant: compare

1747/2 and 810.5

873.5>810.5, consequently , Cu2S is a limiting reactant.

According to equation mole ratio n( Cu2S):n(Cu)= 1:6, then n(Cu)= 6*n(Cu2S)= 6*810.5 = 4863 mol.

m(Cu)= 63.55*4863=

309044 g= 309.044 kg.

% yield = (285 kg/309.044 kg )*100% = 92.2%

Answer: 92.2%